# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 1 Introduction
$\mathbb{O}^{n\times k}=\{P\in\mathbb{R}^{n\times k}\,:\,P^{\operatorname{T}}P=I_{k}\}\subset\mathbb{R}^{n\times k}$ (1.2)
@@
$\mathbb{O}^{n\times k}_{\delta}:=\{P\in\mathbb{R}^{n\times k}\,:\,\|P^{\operatorname{T}}P-I_{k}\|<\delta\}$ (1.3)
@@
Although in general objective function $f$ can be any differentiable function that is well-defined on $\mathbb{O}^{n\times k}_{\delta}$, in practical applications often $f$ is a composition of matrix traces of linear or quadratic forms in $P$. A partial list of most commonly used ones in the literature is given in Table 1, where $A$, $B$, all $A_{i}$, $D$, and all $D_{i}$ are constant matrices, and $A$, $B$, and all $A_{i}$ are at least symmetric and may or may not be positive semidefinite. All but the last two in the table are clearly composed of one or more matrix traces depending on $P$, and the last two are no exceptions! To see that, we notice
@@ 1.1 Review of the NEPv and NPDo Approach
    $H(P)P=P\Omega,\,\,P\in\mathbb{O}^{n\times k}$ (1.5)
    that either is or can be made equivalent to the first order optimality condition, also known as the KKT condition (see section 2 for detail), where $H(P)\in\mathbb{R}^{n\times n}$ is a symmetric matrix-valued function dependent on $P$;
@@
    solve NEPv (1.5) by the self-consistent-field (SCF) iteration: given $P_{0}$,
[…]
    where $H(P_{i-1})\widehat{P}_{i}=\widehat{P}_{i}\Omega_{i}$ is a partial eigendecomposition of $H(P_{i-1})$ for its $k$ largest (or smallest) eigenvalues.
@@
    $\mathscr{H}(P)=P\Lambda,\,\,P\in\mathbb{O}^{n\times k},$ (1.7)
@@
    solve NPDo (1.7) by the self-consistent-field iteration: given $P_{0}$,
[…]
    where $\mathscr{H}(P_{i-1})=\widehat{P}_{i}\Lambda_{i}$ is a polar decomposition [^1] Throughout this paper, a polar decomposition of $B\in\mathbb{R}^{n\times k}$ $(k\leq n)$ refers to $B=P\Omega$ with $P\in\mathbb{O}^{n\times k}$ and $\Omega\in\mathbb{R}^{k\times k}$ being positive semidefinite. The decomposition is unique if and only if $\operatorname{rank}(B)=k$ [38]. The matrix $P$ in the decomposition is called an orthogonal polar factor of $B$. of $\mathscr{H}(P_{i-1})$.
A key prerequisite of the NPDo approach is that, at an optimality $P_{*}$, (1.7) is a polar decomposition of $\mathscr{H}(P_{*})$. This is proved in [66] for SumCT under the condition that all $A_{i}$ are positive semidefinite, and later in this paper, we will prove it for more optimization problems, including those in Table 1 that do not appear in ratio forms. As $\mathscr{H}(P)$ to be decomposed also depends on orthogonal polar factor $P$, we call it a nonlinear polar decomposition with orthogonal polar factor dependency, or NPDo in short. Polar decomposition is often computed via SVD [24] which can be viewed as a special SEP. For that reason, NPDo may also be regarded as a special NEPv.
@@ 1.2 Contributions
    showing the NEPv and NPDo approaches work on each individual atomic function for the approach and, more importantly, any convex composition $\phi\circ T$ of their respective atomic functions, where $\phi(\boldsymbol{x})$ for $\boldsymbol{x}\in\mathbb{R}^{N}$ is convex, each entry of $T(P)\in\mathbb{R}^{N}$ is an atomic function, and the partial derivative of $\phi$ with respect to an entry may be required nonnegative, depending on the particular atomic function that occupies the entry.
@@ 1.3 Organization and Notation
After stating the KKT condition of maximization problem (1.1) in section 2, we divide the rest of this paper into two parts. With maximization problem (1.1) in mind, in Part I we focus on the NPDo approach to solve the KKT condition in three sections: section 3 creates a unifying NPDo framework, including the NPDo Ansatz to guarantee that the KKT condition is an NPDo at optimality of (1.1), the global convergence of the SCF iteration (1.8); section 4 develops a general theory that governs atomic functions for NPDo and show that matrix-trace functions, $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ and $\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ and their powers of order higher than $1$, are atomic functions; finally in section 5, we investigate the NPDo approach for convex compositions $\phi\circ T(P)$ of atomic functions and elaborate on a few $T(P)\in\mathbb{R}^{N}$ of common matrix-trace functions that include those appearing in Table 1. In Part II, we focus on the NEPv approach for the same purpose. It also has three sections to address the corresponding issues: a unifying framework built upon an ansatz – the NEPv Ansatz, the global convergence of the SCF iteration (1.6), atomic functions for NEPv, and their convex compositions $\phi\circ T(P)$ along with a few $T(P)\in\mathbb{R}^{N}$ of common matrix-trace functions. Both frameworks are very similar in appearance, but there are subtle differences in requirements and ease to use, making each have advantages over the other in circumstances. A brief comparison to highlight the major differences between the two approaches is made in section 9. Concluding remarks are drawn in section 10. There are seven appendices at the end to supplement necessary material. Appendix A reviews the canonical angles between subspaces of equal dimensions; appendix B cites a couple of well known inequalities for scalars and establishes a few new ones for matrices to serve the main body of the paper; appendixes C and D contain the proofs of main convergence theorems for NPDo and NEPv, respectively; appendix E briefly outlines the idea to extend our developments to the case under the $M$-inner product, i.e., for the $M$-orthogonal constraint $P^{\operatorname{T}}MP=I_{k}$ instead of $P^{\operatorname{T}}P=I_{k}$; and appendix F refines [67, Theorem 2.2] in the form of the NEPv Ansatz.
@@
    $\Sigma=\operatorname{diag}(\sigma_{1}(B),\sigma_{2}(B),\ldots,\sigma_{s}(B))\in\mathbb{R}^{s\times s},\,\,U\in\mathbb{O}^{m\times s},\,\,V\in\mathbb{O}^{n\times s},$
@@
  • A matrix $A\succ 0\,(\succeq 0)$ means that it is symmetric and positive definite (semi-definite), and accordingly $A\prec 0\,(\preceq 0)$ if $-A\succ 0\,(\succeq 0)$;
  • $\mathbb{O}^{n\times k}$ in (1.2) denotes the Stiefel manifold and $\mathbb{O}^{n\times k}_{\delta}$ in (1.3) is some neighborhood of it; Also frequently, $\mathbb{O}_{D+}^{n\times k}:=\{X\in\mathbb{O}^{n\times k}\,:\,X^{\operatorname{T}}D\succeq 0\}$ given $D\in\mathbb{R}^{n\times k}$.
@@ 2 KKT Condition
Consider maximization problem (1.1) on the Stiefel manifold in its generality. For $P=[p_{ij}]\in\mathbb{O}^{n\times k}_{\delta}$ defined in (1.3), denote by
@@
As an optimization problem on the Stiefel manifold, the first order optimality condition (1.1), also known as the KKT condition, is given by setting the Riemannian gradient of $f$ with respect to the Stiefel manifold $\mathbb{O}^{n\times k}$ at $P$ to zero. It is well-known that the Riemannian gradient of a smooth function $f$ with respect to the Stiefel manifold at $P\in\mathbb{O}^{n\times k}$ can be calculated according to (see, e.g., [1, (3.37)])
$\operatorname{grad}f_{|{{\mathbb{O}}^{n\times k}}}(P)=\Pi_{P}\big{(}\mathscr{H}(P)\big{)}=\mathscr{H}(P)-P\cdot\operatorname{sym}\!\big{(}P^{\operatorname{T}}\mathscr{H}(P)\big{)},$ (2.2)
@@
where the projection $\Pi_{P}(Z):=P-P\operatorname{sym}(P^{\operatorname{T}}Z)$ with $\operatorname{sym}(P^{\operatorname{T}}Z)=(P^{\operatorname{T}}Z+Z^{\operatorname{T}}P)/2$. Setting $\operatorname{grad}f_{|{{\mathbb{O}}^{n\times k}}}(P)=0$ yields the first-order optimality condition:
@@
$\mathscr{H}(P)=P\Lambda\quad\mbox{with}\quad\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\quad P\in\mathbb{O}^{n\times k},$ (2.3)
@@
where $\Lambda=\operatorname{sym}(P^{\operatorname{T}}\mathscr{H}(P))$. The exact form of $\Lambda$, however, is not important, but its symmetry is, for example, it implies that $P^{\operatorname{T}}\mathscr{H}(P)$ is symmetric at any KKT point $P$. The KKT condition (2.3) can also be inferred from treating $P\in\mathbb{O}^{n\times k}$ as the orthogonality constraint $P^{\operatorname{T}}P=I_{k}$ and then using the classical method of Lagrange multipliers for constrained optimization [50]. Geometrically, the condition (2.3) simply asks for that the Euclidean gradient belongs to the normal space to the Stiefel manifold at $P$.
@@
For simple functions, $f(P)=\operatorname{tr}(P^{\operatorname{T}}D)$ or $\operatorname{tr}(P^{\operatorname{T}}AP)$, the KKT condition (2.3) is easily solved. In fact, for the two functions, (2.3) becomes $D=P\Lambda$ or $2AP=P\Lambda$, respectively, which, in consideration of (1.1), tell us that a maximizer can be taken to be an orthogonal polar factor of $D$ [29, 77], or an orthonormal basis matrix of the eigenspace of $A$ associated with its $k$ largest eigenvalues [42, 43, 44, 58], respectively. In both cases, the maximizer as described is considered a close form solution to the respective problem because of the numerical maturity by existing NLA techniques and software [2, 3, 17, 24, 41, 51].
@@
In general, equation (2.3) is not an easy equation to solve in searching for a maximizer with guarantee. For example, in the MAXBET subproblem, simply $f(P)=\operatorname{tr}(P^{\operatorname{T}}D)+\operatorname{tr}(P^{\operatorname{T}}AP)$, the sum of the two simple matrix-trace functions and (2.3) becomes $2AP+D=P\Lambda$ for which there is no existing NLA technique that yields a solution to maximize $f(P)$ with guarantee. Having said that, we point out that the eigenvalue-based method [74], which falls into the NEPv approach, has been demonstrated to be numerically efficient [74] and often produces global maximizers. The MAXBET subproblem is a special case of SumCT. As such, in [66], the NPDo approach has also been successfully applied.
@@
We now formally define the notion of a function being right-unitarily invariant, originally introduced in [46]. It is an important concept that we will frequently refer to in the rest of this paper. However, our definition here differs from [46, Definition 2.1] slightly in that we limit the domain to some neighborhood $\mathbb{O}_{\delta}^{n\times k}$ of $\mathbb{O}^{n\times k}$, rather than the entire space $\mathbb{R}^{n\times k}$ used in [46]. Carefully going through [46], one can see that our definition here is actually sufficient for the development in [46] as it is here.
@@
A function $F\,:\,\mathbb{O}_{\delta}^{n\times k}\to\mathbb{R}^{p\times q}$ is said right-unitarily invariant if
$F(PQ)\equiv F(P)\quad\mbox{for $P\in\mathbb{O}_{\delta}^{n\times k}$ and $Q\in\mathbb{O}^{k\times k}$}.$
@@ 3.1 The NPDo Ansatz
The success of the NPDo approach in [66] rests on a monotonicity lemma which motivates us to formulate the following ansatz to build our framework upon. The key point of the assumption is the ability to generate an improved approximate maximizer $\widetilde{P}$ from a given one $P$, where both the given $P$ and the improved $\widetilde{P}$ may have to come out of possibly a strict subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$. What $\mathbb{P}$ to use depends on the underlying optimization problem (1.1) at hand, as we will repeatedly demonstrate later by concrete examples.
The NPDo Ansatz. Let function $f$ be defined in some neighborhood $\mathbb{O}_{\delta}^{n\times k}$ of $\mathbb{O}^{n\times k}$, and denote by $\mathscr{H}(P)=\frac{\partial f(P)}{\partial P}$. Given $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, if
@@
then there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, where $\omega$ is some positive constant, independent of $P$ and $\widehat{P}$.
For any given $P\in\mathbb{O}^{n\times k}$, by Lemma B.11, there is always $\widehat{P}\in\mathbb{O}^{n\times k}$ such that (3.1) holds with some $\eta>0$, unless for that given $P$, (2.3) holds with $\Lambda\succeq 0$. In fact, we can take $\widehat{P}\in\mathbb{O}^{n\times k}$ to be an orthogonal polar factor of $\mathscr{H}(P)$, which also maximizes $\operatorname{tr}(X^{\operatorname{T}}\mathscr{H}(P))$ over $X\in\mathbb{O}^{n\times k}$ to $\|\mathscr{H}(P)\|_{\operatorname{tr}}$ again by Lemma B.11, in which case $\eta=\|\mathscr{H}(P)\|_{\operatorname{tr}}-\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))$. As far as verifying this ansatz is concerned, it is the desirable aim, $f(\widetilde{P})\geq f(P)+\omega\eta$, that needs to be checked. The necessity of also involving $\mathbb{P}$, a subset of $\mathbb{O}^{n\times k}$, can be best justified by Example 3.1 below.
@@
Consider $f(P)=\operatorname{tr}(P^{\operatorname{T}}AP)+\operatorname{tr}((P^{\operatorname{T}}D)^{2})$ where $0\preceq A\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{n\times k}$. It can be verified that $\mathscr{H}(P)=2AP+2DP^{\operatorname{T}}D$ (see also (4.11) in the next section). Suppose now that (3.1) holds for $P,\,\widehat{P}\in\mathbb{O}^{n\times k}$, or equivalently,
@@
successfully separating $P$ and $\widehat{P}$ from their coupling by $\operatorname{tr}(\widehat{P}^{\operatorname{T}}AP)$. Turning to $\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D)$, we assume that $P^{\operatorname{T}}D\succeq 0$, i.e., $P\in\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$, and let $Q\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ and hence $Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$, implying $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. We get
@@
Combine (3.2), (3.3), and (3.4) to get $f(\widetilde{P})\geq f(P)+\eta$ upon noticing $\operatorname{tr}(\widetilde{P}^{\operatorname{T}}A\widetilde{P})=\operatorname{tr}(\widehat{P}^{\operatorname{T}}A\widehat{P})$. We observe the critical conditions: $P^{\operatorname{T}}D\succeq 0$ and $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$ that ensure (3.4), which we use to separate $P$ and $\widehat{P}$ from their coupling by $\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D)$. The first condition $P^{\operatorname{T}}D\succeq 0$ can be fulfilled by simply starting with $P\in\mathbb{P}$, while the second condition $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$ is made possible by the chosen $Q$ and, as a byproduct, $\widetilde{P}\in\mathbb{P}$, too. Besides this role of making $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$, $Q$ also increases the objective value as a result of the two inequality signs in the derivation of (3.4). In our later use of (3.1), we begin with some $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and then find some $\widehat{P}\in\mathbb{O}^{n\times k}$ such that $\eta>0$ in (3.1), and therefore having a flexibility of judiciously choosing a proper $Q$ becomes a logical necessity.

### v4

@@ 1 Introduction
${\rm St}(k,n)=\{P\in\mathbb{R}^{n\times k}\,:\,P^{\operatorname{T}}P=I_{k}\}\subset\mathbb{R}^{n\times k}$ (1.2)
@@
${\rm St}_{\delta}(k,n):=\{P\in\mathbb{R}^{n\times k}\,:\,\|P^{\operatorname{T}}P-I_{k}\|<\delta\}$ (1.3)
@@
Although in general objective function $f$ can be any differentiable function that is well-defined on ${\rm St}_{\delta}(k,n)$, in practical applications often $f$ is a composition of matrix traces of linear or quadratic forms in $P$. A partial list of most commonly used ones in the literature is given in Table 1, where $A$, $B$, $D$, all $A_{i}$ and $D_{i}$ are constant matrices, and $A$, $B$, and all $A_{i}$ are at least symmetric and may or may not be positive semidefinite. All but the last two in the table are clearly composed of one or more matrix traces depending on $P$, and the last two are no exceptions! To see that, we notice
@@ 1.1 Review of the NEPv and NPDo Approach
    $H(P)P=P\Omega,\,\,P\in{\rm St}(k,n)$ (1.5)
    that either is or can be made equivalent to the first order optimality condition, also known as the KKT condition (see section 2 for detail), where $H(P)\in\mathbb{R}^{n\times n}$ is a symmetric matrix-valued function dependent of $P$;
@@
    solve NEPv (1.5) by the self-consistent-field (SCF) iteration: given $P_{0}$, iteratively
    compute partial eigendecomposition $H(P_{i-1})\widehat{P}_{i}=\widehat{P}_{i}\Omega_{i}$ associated with the $k$ largest (or smallest) eigenvalues of $H(P_{i-1})$ for $\widehat{P}_{i}\in{\rm St}(k,n)$, and postprocess $\widehat{P}_{i}$ to $P_{i}$. (1.6)
@@
    $\mathscr{H}(P)=P\Lambda,\,\,P\in{\rm St}(k,n),$ (1.7)
@@
    solve NPDo (1.7) by the self-consistent-field iteration: given $P_{0}$, iteratively
    compute polar decomposition [^1] Throughout this paper, a polar decomposition of $B\in\mathbb{R}^{n\times k}$ $(k\leq n)$ refers to $B=P\Omega$ with $P\in{\rm St}(k,n)$ and positive semidefinite $\Omega\in\mathbb{R}^{k\times k}$. $\Omega=(B^{\operatorname{T}}B)^{1/2}$ is always unique, but $P\in{\rm St}(k,n)$ is unique if and only if $\operatorname{rank}(B)=k$ [38]. The matrix $P$ in the decomposition is called an orthogonal polar factor of $B$.$\mathscr{H}(P_{i-1})=\widehat{P}_{i}\Lambda_{i}$ of $\mathscr{H}(P_{i-1})$ for $\widehat{P}_{i}\in{\rm St}(k,n)$, and postprocess $\widehat{P}_{i}$ to $P_{i}$. (1.8)
A key prerequisite of the NPDo approach is that, at an optimality $P_{*}$, (1.7) is a polar decomposition of $\mathscr{H}(P_{*})$. This is proved in [66] for SumCT under the condition that all $A_{i}$ are positive semidefinite, and later in this paper, we will prove it for more optimization problems, including those in Table 1 that do not appear in ratio forms. As $\mathscr{H}(P)$ to be decomposed also depends on orthogonal polar factor $P$, we call it a nonlinear polar decomposition with orthogonal polar factor dependency, or NPDo in short. Polar decomposition is often computed via SVD [24] which can be viewed as a special SEP [17]. For that reason, NPDo may also be regarded as a special NEPv.
@@ 1.2 Contributions
    showing the NEPv and NPDo approaches work on each individual atomic function for the approach and, more importantly, any convex composition $\phi\circ T$ of their respective atomic functions, where $\phi(\boldsymbol{x})$ for $\boldsymbol{x}\in\mathfrak{D}\subseteq\mathbb{R}^{N}$ is convex, each entry of $T(P)\in\mathbb{R}^{N}$ is an atomic function, and the partial derivative of $\phi$ with respect to an entry may be required nonnegative, depending on the particular atomic function that occupies the entry.
@@ 1.3 Organization and Notation
After stating the KKT condition of maximization problem (1.1) in section 2, we divide the rest of this paper into two parts. With maximization problem (1.1) in mind, in Part I we focus on the NPDo approach to solve the KKT condition in three sections: section 3 creates a unifying NPDo framework, including the NPDo Ansatz to guarantee that the KKT condition is an NPDo at optimality of (1.1), the global convergence of the SCF iteration (1.8); section 4 develops a general theory that governs atomic functions for NPDo and show that matrix-trace functions, $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ and $\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ and their powers of order higher than $1$, are atomic functions; finally in section 5, we investigate the NPDo approach for convex compositions $\phi\circ T(P)$ of atomic functions and elaborate on a few $T(P)\in\mathbb{R}^{N}$ of common matrix-trace functions that include some of those appearing in Table 1. In Part II, we focus on the NEPv approach for the same purpose. It also has three sections to address the corresponding issues: a unifying framework built upon an ansatz – the NEPv Ansatz, the global convergence of the SCF iteration (1.6), atomic functions for NEPv, and their convex compositions $\phi\circ T(P)$ along with a few $T(P)\in\mathbb{R}^{N}$ of common matrix-trace functions. Both frameworks are very similar in appearance, but there are subtle differences in requirements and ease to use, making each have advantages over the other in circumstances. A brief comparison to highlight the major differences between the two approaches is made in section 9. Concluding remarks are drawn in section 10. There are six appendices at the end to supplement necessary material. Appendix A reviews the canonical angles between subspaces of equal dimensions; appendix B cites a couple of well known inequalities for scalars and establishes a few new ones for matrices to serve the main body of the paper; appendixes C and D contain the proofs of main convergence theorems for NPDo and NEPv, respectively; appendix E briefly outlines the idea to extend our developments to the case under the $M$-inner product, i.e., for the $M$-orthogonal constraint $P^{\operatorname{T}}MP=I_{k}$ instead of $P^{\operatorname{T}}P=I_{k}$; and appendix F refines [67, Theorem 2.2] in the form of the NEPv Ansatz.
@@
  • ${\rm St}(k,n)$ in (1.2) denotes the Stiefel manifold and ${\rm St}_{\delta}(k,n)$ in (1.3) is some neighborhood of it; Also frequently, given $D\in\mathbb{R}^{n\times k}$,
    ${\rm St}(k,n)_{D+}:=\{X\in{\rm St}(k,n)\,:\,X^{\operatorname{T}}D\succeq 0\};$
@@
    $\Sigma=\operatorname{diag}(\sigma_{1}(B),\sigma_{2}(B),\ldots,\sigma_{s}(B))\in\mathbb{R}^{s\times s},\,\,U\in{\rm St}(s,m),\,\,V\in{\rm St}(s,n),$
@@
  • A matrix $A\succ 0\,(\succeq 0)$ means that it is symmetric and positive definite (semi-definite), and accordingly $A\prec 0\,(\preceq 0)$ if $-A\succ 0\,(\succeq 0)$.
@@ 2 KKT Condition
Consider maximization problem (1.1) on the Stiefel manifold ${\rm St}(k,n)$ in its generality. For $P=[p_{ij}]\in{\rm St}_{\delta}(k,n)$ defined in (1.3), denote by
@@
As an optimization problem on the Stiefel manifold, the first order optimality condition (1.1), also known as the KKT condition, is given by setting the Riemannian gradient of $f$ with respect to the Stiefel manifold ${\rm St}(k,n)$ at $P$ to 00. It is well-known that the Riemannian gradient of a smooth function $f$ with respect to the Stiefel manifold at $P\in{\rm St}(k,n)$ can be calculated according to (see, e.g., [1, (3.37)])
$\operatorname{grad}f_{|{{\rm St}(k,n)}}(P)=\Pi_{P}\big{(}\mathscr{H}(P)\big{)}=\mathscr{H}(P)-P\cdot\operatorname{sym}\!\big{(}P^{\operatorname{T}}\mathscr{H}(P)\big{)},$ (2.2)
@@
where the projection $\Pi_{P}(Z):=P-P\operatorname{sym}(P^{\operatorname{T}}Z)$ with $\operatorname{sym}(P^{\operatorname{T}}Z)=(P^{\operatorname{T}}Z+Z^{\operatorname{T}}P)/2$. Setting $\operatorname{grad}f_{|{{\rm St}(k,n)}}(P)=0$ yields the first-order optimality condition:
@@
$\mathscr{H}(P)=P\Lambda\quad\mbox{with}\quad\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\quad P\in{\rm St}(k,n),$ (2.3)
@@
where $\Lambda=\operatorname{sym}(P^{\operatorname{T}}\mathscr{H}(P))$. The exact form of $\Lambda$, however, is not important, but its symmetry is, for example, it implies that $P^{\operatorname{T}}\mathscr{H}(P)$ is symmetric at any KKT point $P$. The KKT condition (2.3) can also be inferred from treating $P\in{\rm St}(k,n)$ as the orthogonality constraint $P^{\operatorname{T}}P=I_{k}$ and then using the classical method of Lagrange multipliers for constrained optimization [50]. Geometrically, the condition (2.3) simply asks for that the Euclidean gradient belongs to the normal space to the Stiefel manifold at $P$.
@@
For simple functions, $f(P)=\operatorname{tr}(P^{\operatorname{T}}D)$ or $\operatorname{tr}(P^{\operatorname{T}}AP)$, the KKT condition (2.3) can be considered as solved. In fact, for the two functions, (2.3) becomes $D=P\Lambda$ or $2AP=P\Lambda$, respectively, which, in consideration of (1.1), tell us that a maximizer can be taken to be an orthogonal polar factor of $D$ [29, 77], or an orthonormal basis matrix of the eigenspace of $A$ associated with its $k$ largest eigenvalues [42, 43, 44, 58], respectively. In both cases, the maximizer as described is considered a close form solution to the respective problem because of the numerical maturity by existing NLA techniques and software [2, 3, 17, 24, 41, 51].
@@
In general, equation (2.3) is not an easy equation to solve in searching for a maximizer of (1.1) with guarantee. For example, in the MAXBET subproblem, simply $f(P)=\operatorname{tr}(P^{\operatorname{T}}D)+\operatorname{tr}(P^{\operatorname{T}}AP)$, the sum of the two simple matrix-trace functions and (2.3) becomes $2AP+D=P\Lambda$ for which there is no existing NLA technique that yields a solution to maximize $f(P)$ with guarantee. Having said that, we point out that the eigenvalue-based method [74], which falls into the NEPv approach, has been demonstrated to be numerically efficient [74] and often produces global maximizers. The MAXBET subproblem is a special case of SumCT. As such, in [66], the NPDo approach has also been successfully applied.
@@
We now formally define the notion of a function being right-unitarily invariant, originally introduced in [46]. It is an important concept that we will frequently refer to in the rest of this paper. However, our definition here differs from [46, Definition 2.1] slightly in that we limit the domain to some neighborhood ${\rm St}_{\delta}(k,n)$ of the Stiefel manifold ${\rm St}(k,n)$, rather than the entire space $\mathbb{R}^{n\times k}$ used in [46]. Carefully going through [46], one can see that our definition here is actually sufficient for the development in [46] as it is here.
@@
A function $F\,:\,{\rm St}_{\delta}(k,n)\to\mathbb{R}^{p\times q}$ is said right-unitarily invariant if
$F(PQ)\equiv F(P)\quad\mbox{for $P\in{\rm St}_{\delta}(k,n)$ and $Q\in{\rm St}(k,k)$}.$
@@ 3.1 The NPDo Ansatz
The success of the NPDo approach in [66] rests on a monotonicity lemma which motivates us to formulate the following ansatz to build our framework upon. The key point of the assumption is the ability to generate an improved approximate maximizer $\widetilde{P}$ from a given one $P$, where both the given $P$ and the improved $\widetilde{P}$ may have to come out of possibly a strict subset $\mathbb{P}$ of ${\rm St}(k,n)$. What $\mathbb{P}$ to use depends on the underlying optimization problem (1.1) at hand, as we will repeatedly demonstrate later by concrete examples.
The NPDo Ansatz. Let function $f$ be defined in some neighborhood ${\rm St}_{\delta}(k,n)$ of ${\rm St}(k,n)$, and denote by $\mathscr{H}(P)=\frac{\partial f(P)}{\partial P}$. Given $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and $\widehat{P}\in{\rm St}(k,n)$, if
@@
then there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, where $\omega$ is some positive constant, independent of $P$ and $\widehat{P}$.
For any given $P\in{\rm St}(k,n)$, by Lemma B.11, there is always $\widehat{P}\in{\rm St}(k,n)$ such that (3.1) holds with some $\eta>0$, unless for that given $P$, (2.3) holds with $\Lambda\succeq 0$. In fact, we can take $\widehat{P}\in{\rm St}(k,n)$ to be an orthogonal polar factor of $\mathscr{H}(P)$, which also maximizes $\operatorname{tr}(X^{\operatorname{T}}\mathscr{H}(P))$ over $X\in{\rm St}(k,n)$ to $\|\mathscr{H}(P)\|_{\operatorname{tr}}$ again by Lemma B.11, in which case $\eta=\|\mathscr{H}(P)\|_{\operatorname{tr}}-\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))$. Hence, for the purpose of solving (1.1), we may relax the ansatz to $\eta\geq 0$ only. As far as verifying this ansatz is concerned, it is the desirable aim, $f(\widetilde{P})\geq f(P)+\omega\eta$, that needs to be checked. The necessity of also involving $\mathbb{P}$, a subset of ${\rm St}(k,n)$, can be best justified by Example 3.1 below.
@@
Consider $f(P)=\operatorname{tr}(P^{\operatorname{T}}AP)+\operatorname{tr}((P^{\operatorname{T}}D)^{2})$ where $0\preceq A\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{n\times k}$. It can be verified that $\mathscr{H}(P)=2AP+2DP^{\operatorname{T}}D$ (see also (4.11) in the next section). Suppose now that (3.1) holds for $P,\,\widehat{P}\in{\rm St}(k,n)$, or equivalently,
@@
successfully separating $P$ and $\widehat{P}$ from their coupling by $\operatorname{tr}(\widehat{P}^{\operatorname{T}}AP)$. Turning to $\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D)$, we assume that $P^{\operatorname{T}}D\succeq 0$, i.e., $P\in\mathbb{P}={\rm St}(k,n)_{D+}$, and let $Q\in{\rm St}(k,k)$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ and hence $Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$, implying $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. We get
@@
Combine (3.2), (3.3), and (3.4) to get $f(\widetilde{P})\geq f(P)+\eta$ upon noticing $\operatorname{tr}(\widetilde{P}^{\operatorname{T}}A\widetilde{P})=\operatorname{tr}(\widehat{P}^{\operatorname{T}}A\widehat{P})$. We observe the critical conditions: $P^{\operatorname{T}}D\succeq 0$ and $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$ that ensure (3.4), which we use to separate $P$ and $\widehat{P}$ from their coupling by $\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D)$. The first condition $P^{\operatorname{T}}D\succeq 0$ can be fulfilled by simply starting with $P\in\mathbb{P}$, while the second condition $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$ is made possible by the chosen $Q$ and, as a byproduct, $\widetilde{P}\in\mathbb{P}$, too. Besides this role of making $Q^{\operatorname{T}}\widehat{P}^{\operatorname{T}}D\succeq 0$, $Q$ also increases the objective value as a result of the two inequality signs in the derivation of (3.4). In our later use of (3.1), we begin with some $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and then find some $\widehat{P}\in{\rm St}(k,n)$ such that $\eta>0$ in (3.1), and therefore having a flexibility of judiciously choosing a proper $Q$ becomes a logical necessity.